# Constraining Power of Wavelet vs. Power Spectrum Statistics for CMB Lensing and Weak Lensing with Learned Binning

### v1

@@ 5.2 CMB Lensing X Weak Lensing
ACT $\times$ Euclid DR2 3.4 / 3.9 1.3 / 1.6
Planck $\times$ Euclid DR2 2.7 / 3.1 1.0 / 1.3
SPT $\times$ Euclid DR2 2.3 / 2.8 0.7 / 0.9
SO $\times$ Euclid DR2 2.6 / 3.1 0.7 / 0.9
@@
ACT $\times$ Euclid DR2 4.0 / 4.5 3.5 / 4.0
Planck $\times$ Euclid DR2 3.3 / 3.7 2.7 / 3.0
SPT $\times$ Euclid DR2 2.4 / 2.9 1.5 / 1.6
SO $\times$ Euclid DR2 2.7 / 3.2 1.4 / 1.6
@@
For cross-$C_{\ell}$’s, we have two spectra coming from our two tomographic bins for WL, leading to 6140 $C_{\ell}$’s. We order each power spectrum based on $\ell$ and then append them. We lock a pivot at the transition between tomographic bins. We find improvements by adding two more arbitrary pivots, leading to a $5$ dimensional search space for the pivots, but no significant improvements going beyond this.
For WPH, for each tomographic bin we have $S_{00}$, $S_{01}$, $S_{10}$, $S_{11}$, $C_{01}$, and $C_{10}$ coefficients. For $C_{01}$ and $C_{10}$, we also have four different opening angles that we split into different subclasses of statistics. From this, for each of the two tomographic bins we have 12 subclasses within the WPH summary statistic, leading to a total of 24 subclasses. Within each subclass, we order based on $j$. For $C_{01}$ ($C_{10}$), we order based on $j_{2}$ ($j_{1}$). When multiple coefficients have the same opening angle and $j_{2}$ ($j_{1})$ value, we order based on $j_{1}$ ($j_{2}$). We lock pivots at the transitions between each subclass of the WPH statistics. Since this configuration already has a relatively large number of degrees of freedom for binning, we do not add any additional pivots.
As an example of the power of the learned binning, we can compare constraints with the learned binning compared to an initial binning. As the simplest possible initial binning to compare to, we look at linear binning with no pivots. Constraints are shown in Figure 6 for Planck WPH. Constraints improve by factors of $2.7$ for $\Omega_{m}$ and $2.3$ for $\sigma_{8}$ when using a learned binning. We verify the convergence of this method in Appendix B.
@@
All of these results show that WPH is able to extract information from non-Gaussianities in the lensing. These gains are substantially larger than those found for CMB lensing alone, which is expected: weak-lensing – CMB-lensing cross-correlations peak at significantly lower redshift ($z\sim 1$) compared to CMB lensing auto-correlations ($z\sim 2$–6). At these lower redshifts the matter distribution is more evolved and inherently more non-Gaussian, providing WPH with additional information to extract beyond that available to power spectra alone. These findings are consistent with those of previous WL applications in the literature (e.g. Cheng et al., 2020; Gatti et al., 2023; Cheng et al., 2025; Prat et al., 2025; Jeffrey et al., 2025).
@@
To test the convergence of our results, we repeat this entire process (including finding pivots), while lowering the total number of derivatives used. We check how the constraints vary as a function of the number of derivative realizations used and find that the results have sufficiently converged. All results are shown in Appendix B.
6 Conclusion
@@
In this work, we study CMB lensing convergence and galaxy weak lensing convergence data with a variety of summary statistics. To reduce the dimensionality of our summary statistics, we introduce a novel, learned-binning approach. Like other data reduction methods such as MOPED (Heavens et al., 2000), our algorithm attempts to maximize Fisher information in the compressed statistic. Since our method depends on fewer parameters, it is less susceptible to noise arising from a limited number of simulations. Our method also never requires the inversion of the original data covariance matrix, something that becomes impossible if the number of simulations is less than the length of the summary statistic. By construction, our learned binning approach is robust to overfitting and can improve constraints compared to linear binning by factors up to $2.7$ for the surveys and observables considered in this work.
@@
For CMB lensing convergence data, we compare the higher-order statistic of the wavelet scattering transform to the standard angular power spectrum. We find that these summary statistics lead to similar constraints for SPT and SO, while WST outperforms for Planck and ACT. When we cross our CMB lensing simulated data with Euclid DR2 galaxy weak lensing convergence, we find that the higher-order statistic of the wavelet phase harmonics significantly outperforms the cross power spectrum. Constraints are improved by factors of up to $3.8$. Compared to constraints from CMB lensing convergence data alone, the cross-survey constraints exhibit stronger degeneracies. All contour plots are presented in Appendix A.
@@
With WPH showing the most promise of the two higher-order statistics, further generalizations of WPH could prove useful. One potential next step would be investigating the use of WPH statistics besides $S_{00}$, $S_{01}$, $S_{10}$, $S_{11}$, $C_{01}$, and $C_{10}$. WPH and WST also require us to break the sky into patches, which limits the scales that we are able to probe. Generalizing our pipeline to spherical wavelets (McEwen et al., 2021) could circumvent this issue. Spherical wavelets have been previously used for CMB analysis (McEwen et al., 2007; Surrao & Hill, 2024).
@@
We also provide a novel method for the comparison of multiple summary statistics without sacrificing interpretability, which will become increasingly more important as new summary statistics are constantly being introduced.
A physical interpretation of these results follows from the discussion in the previous section. WL-CMBL cross-correlations peak at lower redshifts ($z\sim 1$) than CMBL convergence auto-correlations ($z\sim 2$–6), where the matter distribution is more evolved and therefore more non-Gaussian. This makes non-Gaussian statistics particularly powerful in the cross-correlation case. While applications of non-Gaussian statistics to weak lensing alone are already well established (e.g. Cheng et al., 2020; Gatti et al., 2023; Cheng et al., 2025; Chudaykin et al., 2025; Sugiyama et al., 2025; Prat et al., 2025; Jeffrey et al., 2025), their use in WL–CMBL convergence cross-correlation data has not been explored before. One reason is that it requires forward modeling of the CMB lensing reconstruction process, including all relevant observational systematics, which is highly non-trivial. Nonetheless, the results presented here suggest that such an effort would be worthwhile.
7 Acknowledgements
@@ Appendix A Constraints for All Surveys
Appendix B Convergence for All Surveys
We test the convergence of our methodology by reducing the number of derivatives available and checking how the contours change. The number of realizations for generating both $\Omega_{m}$ and $\sigma_{8}$ derivatives are changed simultaneously. For brevity, we only show the convergence contours for $\kappa_{\text{CMB}}\times\kappa_{\text{WL}}$ in Figures B.1 and B.2. Constraints for auto $\kappa_{\text{CMB}}$ all converge to similar levels. Additionally, we show how $1\sigma$ constraints on $\Omega_{m}$ and $\sigma_{8}$ converge individually for cross-$C_{\ell}$’s and WPH, in Figures B.3 and B.4, respectively.
@@ Appendix B Convergence for All Surveys
Figure B.3: Convergence of $\Omega_{m}$ (left) and $\sigma_{8}$ (right) constraints as a function of the number of derivative realizations used to evaluate cross-$C_{\ell}$’s from $\kappa_{\text{CMB}}\times\kappa_{\text{WL}}$. By the time 50 realizations are used, constraints have converged to within $\sim 30\%$ of their final values. The high degeneracy of the parameters leads to the two plots looking nearly identical.
@@
Figure B.4: Same as Figure B.3, but using WPH as the summary statistic. Even when only $25$ realizations are used to evaluate derivatives, the results have converged to within $10\%$ of their final values.
Appendix C Binning Matrix Construction
@@
In order to properly construct a binning matrix, the function $f(x):[1,M]\rightarrow[1,N]$ defined in Section 4 must satisfy certain conditions. Since the first entry of the summary statistic is mapped to the first bin, we must have $f(1)=1$. Similarly, requiring that the last entry of the summary statistic is mapped to the last bin gives the requirement $f(M)=N$. Bins should consist of entries in the summary statistic which are all adjacent to one another (no gaps in the bins), which demands that $f(x)$ is a non-decreasing function. Finally, we want to ensure that we have no empty bins. There are numerous constraints one could impose on $f(x)$ to guarantee this property, and in this work we choose to require $f^{\prime}(x)\leq 1$.
@@
To construct a binning matrix, we would generally have the $m^{\text{th}}$ column of the binning matrix be nonzero (and equal to $1$) only at the integer index closest to $f(m)$. In this work, we slightly generalize binning to allow for one summary statistic entry to be mapped into two adjacent bins. For example, if $f(2)=3.4$, the $2^{\text{nd}}$ summary statistic entry would be divided $60\%$ into bin $3$ and $40\%$ into bin $4$. This splitting ensures that the determinant of the Fisher matrix is a continuous function with respect to pivot positions.
Appendix D Gaussianity of the Likelihood
@@
The Fisher forecast methods we use rely on our summary statistics following a Gaussian distribution. With enough binning, the central limit theorem will ensure that our statistics are indeed Gaussian. We test whether this is the case by producing a $\chi^{2}$ distribution of the statistics as in Valogiannis et al. (2024a). Specifically, if a summary statistic on any given realization is given by $\text{\bf O}_{i}$ with an average value of $\bar{\text{\bf O}}$ and covariance matrix $C$, the $\chi^{2}$ values can be given by:
$\chi^{2}=(\text{\bf O}_{i}-\bar{\text{\bf O}})^{T}C^{-1}(\text{\bf O}_{i}-\bar{\text{\bf O}}).$ (D1)
@@
We perform this test with our binned data. Specifically, we use $500$ fiducial realizations in learning the binning, and then apply that binning to the other $500$ realizations which we use to get $\chi^{2}$ values. Since we use $15$ bins, if our statistics are Gaussian distributed, the calculated $\chi^{2}$ values should follow the theoretical $\chi^{2}$ distribution with $15$ degrees of freedom. We verify this in Figure D.1 where we find no significant differences between the $\chi^{2}$ distribution of our data, the theoretical $\chi^{2}$ distribution, and a Gaussian distribution from samples randomly generated with the same mean and covariance as our data.
Refer to caption Refer to caption
Refer to caption Refer to caption
@@
Figure D.1: Probability density functions of the $\chi^{2}$ distribution for all summary statistics used. One survey is picked at random for each summary statistic for brevity, but results for other surveys are comparable. $500$ fiducial realizations (orange) are used for the distributions after being binned with a binning learned by the other $500$ fiducial simulations. Alongside the distribution of the data, we show the theoretical $\chi^{2}$ distribution with $15$ degrees of freedom (black) and samples drawn from a Gaussian distribution with the same mean and covariance as the data (blue). We find no evidence of significant departures from Gaussianity for any of our statistics.

### v2

@@ V.2 CMB Lensing X Weak Lensing
ACT $\times$ Euclid DR2 4.09 / 4.97 1.52 / 1.96
Planck $\times$ Euclid DR2 2.79 / 3.39 1.20 / 1.56
SPT $\times$ Euclid DR2 2.57 / 3.29 0.76 / 0.95
SO $\times$ Euclid DR2 2.67 / 3.39 0.86 / 1.10
@@
ACT $\times$ Euclid DR2 3.88 / 4.86 4.07 / 4.84
Planck $\times$ Euclid DR2 2.89 / 3.51 2.80 / 3.30
SPT $\times$ Euclid DR2 2.41 / 3.08 1.68 / 1.96
SO $\times$ Euclid DR2 2.56 / 3.24 1.80 / 2.16
@@
For cross-$C_{\ell}$’s, we have two spectra coming from our two tomographic bins for WL, leading to 6140 $C_{\ell}$’s. We order each power spectrum based on $\ell$ and then append them. We lock a pivot at the transition between tomographic bins and add two more arbitrary pivots so that each tomographic bin can get one pivot like the auto $C_{\ell}$’s of the previous section. This leads to a $5$-dimensional search space for the pivots, and we find no significant improvements beyond this. Across surveys, we find that the first $\sim 500$ $C_{\ell}$’s from the higher redshift tomographic bin are the most heavily weighted.
For WPH, for each tomographic bin we have $S_{00}$, $S_{01}$, $S_{10}$, $S_{11}$, $C_{01}$, and $C_{10}$ coefficients. For $C_{01}$ and $C_{10}$, we also have four different opening angles that we split into different subclasses of statistics. From this, for each of the two tomographic bins, we have 12 subclasses within the WPH summary statistic, leading to a total of 24 subclasses. Within each subclass, we order based on $j$. For $C_{01}$ ($C_{10}$), we order based on $j_{2}$ ($j_{1}$). When multiple coefficients have the same opening angle and $j_{2}$ ($j_{1})$ value, we order based on $j_{1}$ ($j_{2}$). We lock pivots at the transitions between each subclass of the WPH statistics. Since this configuration already has a relatively large number of degrees of freedom for binning, we do not add any additional pivots. Across surveys, $S_{00}$ and $S_{11}$ consistently receive the largest number of bins.
As an example of the power of the learned binning, we can compare constraints with the learned binning compared to an initial binning. As the simplest possible initial binning to compare to, we look at linear binning with no pivots. Constraints are shown in Figure 6 for ACT WPH. Constraints improve by factors of $2.7$ for $\Omega_{m}$ and $2.5$ for $\sigma_{8}$ when using a learned binning. We verify the convergence of this method in Appendix B.
@@
Our results show that WPH can extract information from non-Gaussianities in the lensing. Compared to the lack of gains found for CMB lensing alone, this is expected: weak-lensing – CMB-lensing cross-correlations peak at significantly lower redshift ($z\sim 1$) compared to CMB lensing auto-correlations ($z\sim 2$–6). At these lower redshifts, the matter distribution is more evolved and inherently more non-Gaussian, providing WPH with additional information to extract beyond that available to power spectra alone. These findings are consistent with those of previous WL applications in the literature [e.g. 11, 18, 9, 39, 25].
@@
To test the convergence of our results, we repeat this entire process (including finding pivots) while lowering the total number of derivatives used. We examine how the constraints vary with the number of derivative realizations used and find that the results have sufficiently converged. All results are shown in Appendix B.
VI Conclusion
@@
In this work, we study CMB lensing convergence and galaxy weak-lensing convergence data using a variety of summary statistics. To reduce the dimensionality of our summary statistics, we introduce a novel, learned-binning approach. Like other data reduction methods such as MOPED [22], our algorithm attempts to maximize Fisher information in the compressed statistic. Since our method uses fewer parameters, it is less susceptible to noise from a limited number of simulations. Our method also never requires inverting the original data covariance matrix, which becomes impossible if the number of simulations is less than the length of the summary statistic. By construction, our learned binning approach is robust to overfitting and can improve constraints relative to linear binning by up to a factor of 2.7 for the surveys and observables considered in this work.
@@
For CMB lensing convergence data, we compare the higher-order statistics of the wavelet scattering transform to the standard angular power spectrum. We find that these summary statistics lead to similar constraints for all of the surveys considered. When we cross our CMB lensing simulated data with Euclid DR2 galaxy weak lensing convergence, we find that the higher-order statistic of the wavelet phase harmonics significantly outperforms the cross power spectrum. Constraints are improved by factors of up to $3.4$. Compared to constraints from CMB lensing convergence data alone, the cross-survey constraints exhibit stronger degeneracies. All contour plots are presented in Appendix A.
@@
With WPH showing the most promise of the two higher-order statistics, further generalizations of WPH could prove useful. One potential next step would be investigating the use of WPH statistics besides $S_{00}$, $S_{01}$, $S_{10}$, $S_{11}$, $C_{01}$, and $C_{10}$. WPH and WST also require us to break the sky into patches, which limits the scales we can probe. Generalizing our pipeline to spherical wavelets [32] could circumvent this issue. Spherical wavelets have been previously used for CMB analysis [31, 44].
@@
We also provide a novel method for comparing multiple summary statistics without sacrificing interpretability, which will become increasingly important as new summary statistics are introduced.
A physical interpretation of these results follows from the discussion in the previous section. WL-CMBL cross-correlations peak at lower redshifts ($z\sim 1$) than CMBL convergence auto-correlations ($z\sim 2$–6), where the matter distribution is more evolved and therefore more non-Gaussian. This makes non-Gaussian statistics particularly powerful in the cross-correlation case. While applications of non-Gaussian statistics to weak lensing alone are already well established [e.g. 11, 18, 9, 12, 43, 39, 25], their use in WL–CMBL convergence cross-correlation data has not been explored before. One reason is that it requires forward modeling of the CMB lensing reconstruction process, including all relevant observational systematics, which is highly non-trivial. Nonetheless, the results presented here suggest that such an effort would be worthwhile.
VII Acknowledgements
@@ Appendix A Constraints for All Surveys
Appendix B Contour Convergence for All Surveys
We test the convergence of our methodology by reducing the number of available derivatives and observing how the contours change. The number of realizations for generating both $\Omega_{m}$ and $\sigma_{8}$ derivatives is changed simultaneously. For brevity, we only show the convergence contours for $\kappa_{\text{CMB}}\times\kappa_{\text{WL}}$ statistics in Figures B.1 and B.2. Constraints for auto $\kappa_{\text{CMB}}$ statistics all converge to similar levels. Additionally, we show how $1\sigma$ constraints on $\Omega_{m}$ and $\sigma_{8}$ converge individually for cross-$C_{\ell}$’s and WPH, in Figures B.3 and B.4, respectively.
@@ Appendix B Contour Convergence for All Surveys
Figure B.3: Convergence of $\Omega_{m}$ (left) and $\sigma_{8}$ (right) constraints as a function of the number of derivative realizations used to evaluate cross-$C_{\ell}$’s from $\kappa_{\text{CMB}}\times\kappa_{\text{WL}}$. By the time 50 realizations are used, constraints have converged to within $\sim 15\%$ of their final values. The high degeneracy of the parameters makes the two plots look nearly identical.
@@
Figure B.4: Same as Figure B.3, but using WPH as the summary statistic. By the time 50 realizations are used, constraints have converged to within $\sim 5\%$ of their final values.
Appendix C Derivative Convergence
This section focuses only on the convergence of derivatives of our summary statistics with respect to $\Omega_{m}$. This is done for the sake of brevity; $\sigma_{8}$ derivatives converge to similar levels. We have sets of maps for three different values of $\Omega_{m}$: the fiducial value and the fiducial value $\pm\Delta\Omega_{m}$ with $\Delta\Omega_{m}=0.01$. In this work, for any summary statistic $S$, we calculate our derivatives as:
$\frac{dS}{d\Omega_{m}}=\frac{\langle S(\Omega_{m,fid}+\Delta\Omega_{m})-S(\Omega_{m,fid}-\Delta\Omega_{m})\rangle}{2\Delta\Omega_{m}},$ (26)
where the average is done over the $100$ realizations of these maps. Each pair of maps with $\pm\Delta\Omega_{m}$ shares the same initial seeding, leading to low shot-noise levels due to cancellation. We will denote this derivative estimate as $S^{\prime}$.
Alternatively, we could have defined the derivative as:
$\frac{dS}{d\Omega_{m}}=\frac{\langle S(\Omega_{m,fid})\rangle-\langle S(\Omega_{m,fid}-\Delta\Omega_{m})\rangle}{\Delta\Omega_{m}},$ (27)
where the averages are now over each individual set of realizations, since these maps do not share the initial seed. We will denote this derivative estimate as $\tilde{S}^{\prime}$.
[…]
$\Delta\Omega_{m}(S^{\prime}-\tilde{S}^{\prime})/(\sigma(S)/\sqrt{100}),$ (28)
where $\sigma(S)$ is the elementwise standard deviation of the summary statistic estimated from our fiducial realizations. The factor of $\sqrt{100}$ accounts for the majority of the noise coming from $\langle S(\Omega_{m,fid}-\Delta\Omega_{m})\rangle$, which is an average of $100$ realizations. If the summary statistics have no cross correlations, Eq. 28 values should follow a Gaussian distribution. To account for cross correlations, we sample $100$ vectors from the covariance matrix of the fiducial realizations with mean zero and average them elementwise. We will call this averaged sample vector $D$ and plot out histograms of:
$D/(\sigma(S)/\sqrt{100}).$ (29)
If our derivative estimates are consistent with each other, the distributions from Eq. 28 and Eq. 29 should match. We show examples of the absolute values of these histograms with $68\%$ error bars on the sampled histogram in Figure C.1. We find no significant differences between histograms for any summary statistic across surveys, so we conclude that our derivative estimates have sufficiently converged.
Figure C.1: Distributions of normalized derivative discrepancies with $68\%$ error bars on histograms sampled from the fiducial covariance. One survey is picked at random for each summary statistic for brevity, but results for other surveys are comparable. Additionally, only $\Omega_{m}$ derivatives are shown for brevity, but $\sigma_{8}$ derivatives converge to comparable levels.
Appendix D Learned Binning Details
@@
In order to properly construct a binning matrix, the function $f(x):[1,M]\rightarrow[1,N]$ defined in Section IV must satisfy certain conditions. Since the first entry of the summary statistic is mapped to the first bin, we must have $f(1)=1$. Similarly, requiring that the last entry of the summary statistic is mapped to the last bin gives the requirement $f(M)=N$. Bins should consist of entries in the summary statistic that are adjacent to one another (no gaps in the bins), which requires that $f(x)$ be a non-decreasing function. Finally, we want to ensure there are no empty bins. There are numerous constraints one could impose on $f(x)$ to guarantee this property, and in this work we choose to require $f^{\prime}(x)\leq 1$.
@@
To construct a binning matrix, we would generally have the $m^{\text{th}}$ column of the binning matrix be nonzero (and equal to $1$) only at the integer index closest to $f(m)$. In this work, we slightly generalize binning to allow for one summary statistic entry to be mapped into two adjacent bins. For example, if $f(2)=3.4$, the $2^{\text{nd}}$ summary statistic entry would be divided $60\%$ into bin $3$ and $40\%$ into bin $4$. This splitting ensures that the determinant of the Fisher matrix is a continuous function of pivot positions.
As mentioned in Section IV, our function $f$ is a linear interpolation between pivots, so that we perform optimization in a low-dimensional space. Once we have picked a number of pivots, our loss function is the negative logarithm of the determinant of the Fisher matrix. We include a prior term in our loss function to ensure that the pivots satisfy the conditions explained earlier in this section. We optimize this loss function using the differential evolution method in SciPy [51] with a population size of $20$. The population is initially randomly seeded throughout the pivot space.
Appendix E Gaussianity of the Likelihood
@@
The Fisher forecast methods we use rely on our summary statistics, which follow a Gaussian distribution. With enough binning, the central limit theorem will ensure that our statistics are indeed Gaussian. We test whether this is the case by producing a $\chi^{2}$ distribution of the statistics as in Valogiannis et al. [48]. Specifically, if a summary statistic on any given realization is given by $\text{\bf O}_{i}$ with an average value of $\bar{\text{\bf O}}$ and covariance matrix $C$, the $\chi^{2}$ values can be given by:
$\chi^{2}=(\text{\bf O}_{i}-\bar{\text{\bf O}})^{T}C^{-1}(\text{\bf O}_{i}-\bar{\text{\bf O}}).$ (30)
@@
We perform this test with our binned data. Specifically, we use $500$ fiducial realizations to learn the binning, then apply that binning to the remaining $500$ realizations to obtain $\chi^{2}$ values. Since we use $15$ bins, if our statistics are Gaussian distributed, the calculated $\chi^{2}$ values should follow the theoretical $\chi^{2}$ distribution with $15$ degrees of freedom. We verify this in Figure E.1 where we find no significant differences between the $\chi^{2}$ distribution of our data, the theoretical $\chi^{2}$ distribution, and a Gaussian distribution from samples randomly generated with the same mean and covariance as our data.
@@
Figure E.1: Probability density functions of the $\chi^{2}$ distribution for all summary statistics used. One survey is picked at random for each summary statistic for brevity, but results for other surveys are comparable. $500$ fiducial realizations (orange) are used for the distributions after being binned with a binning learned by the other $500$ fiducial simulations. Alongside the data distribution, we show the theoretical $\chi^{2}$ distribution with $15$ degrees of freedom (black) and samples drawn from a Gaussian distribution with the same mean and covariance as the data (blue). We find no evidence of significant departures from Gaussianity for any of our statistics.